Mynarc Corporation produces fertilizer and distributes the product by using his tanker trucks. Mynarc

uses budgeted fleet hours to allocate variable manufacturing overhead. The following information pertains to the company's manufacturing overhead data:

Budgeted output units                                                                                                675 truckloads

Budgeted fleet hours                                                                                                             540 hours

Budgeted variable manufacturing overhead costs for 675 loads                                    $82,350

Actual output units produced and delivered                                                        630 truckloads

Actual fleet hours                                                                                                                  436 hours

Actual variable manufacturing overhead costs                                                                 $77,490

1a. What is the flexible-budget amount for variable manufacturing overhead and What is the flexible-budget variance for variable manufacturing overhead?
Answer #1
Budgeted variable manufacturing per unit 122 =82350/675
Flexible-budget amount for variable manufacturing overhead 76860 =630*122
Flexible-budget variance for variable manufacturing overhead 630 Unfavorable =77490-76860